Be sure to answer all parts.

How many total moles of ions are released when the following sample dissolves completely in water? Enter your answer in scientific notation.

7.72 × 10^-4 g of Ca(NO3)2: ___ × 10^_____mol
1
Expert's answer
2020-12-06T13:37:44-0500

Ca(NO3)2 "\rightarrow" Ca2+ + 2NO3-

moles of Ca(NO3)2 = "\frac{7.72× 10^{-4}}{164.08}"


moles of Ca(NO3)2 = 4.7× 10-6 moles

Total number of ions = 3

so according to stoichiometry total concentration of ions = 3 × 4.7× 10-6 moles

which is equal to 14.1× 10-6 moles